OpenStudy (anonymous):

A=4i+4j, B=5i -2j Calculate the angle between the vectors A and B (i have an answer but is not the same as my teachers answer (not even in the ball park) just want to see what you guys get...

OpenStudy (anonymous):

∼66.8 degrees

OpenStudy (anonymous):

\[\frac{A•B}{||A||.||B||} = \cos(\Theta)\] Where • is the scalar product (or dot product) between 2 vectors. Appling this on your problem, we've got: \[\frac{20+(-8)}{\sqrt{32}*\sqrt{29}} = \cos(\Theta)\] using acos on both sides: \[acos{\frac{12}{\sqrt{928}}} = \Theta \approx 66,80º\]